Consider the surface given by the parametric equations r=< u^2,2ucos(v),3usin(v)> . Let P be the point (4,0,6). Find an equation of the tangent plane to the surface at the point P.
Answer #1

r={u²,2ucosv,3usinv) x=u² ,y=2ucosv ,z=3usinv Sinu (*)+(5)*=ucos®o+usins *(*)**{*}*=+[cosºu+sin?o} +x=-=0 + f(x,y,z)=0 20=4Eqn of tangent plane Әf, Əf Əf, əx(-x)+ + (z-7.)=0 (2–4)+0(y=0)-1}{3–6)=0 (2–4)-(2-6)=0 3x-12–42+24=0 3x-4x+12=0